# Fast, Accurate and Perturbative Forward Modeling of Galaxy Clustering Part II: Redshift Space

### v1

@@ Abstract
Forward modeling the galaxy density within the Effective Field Theory of Large Scale Structure (EFT of LSS) enables field-level analyses that are robust to theoretical uncertainties. At the same time, they can maximize the constraining power from galaxy clustering on the scales amenable to perturbation theory. In order to apply the method to galaxy surveys, the forward model must account for the full observational complexity of the data. In this context, a major challenge is the inclusion of redshift space distortions (RSDs) from the peculiar motion of galaxies. Here, we present improvements in the efficiency and accuracy of the RSD modeling in the perturbative LEFTfield forward model. We perform a detailed quantification of the perturbative and numerical error for the prediction of momentum, velocity and the redshift-space matter density. Further, we test the recovery of cosmological parameters at the field level, namely the growth rate $f$, from simulated halos in redshift space. For a rigorous test and to scan through a wide range of analysis choices, we fix the linear (initial) density field to the known ground truth but marginalize over all unknown bias coefficients and noise amplitudes. With a third-order model for gravity and bias, our results yield $<1\,\%$ statistical and $<1.5\,\%$ systematic error. The computational cost of the redshift-space forward model is only $\sim 1.5$ times of the rest frame equivalent, enabling future field-level inference that simultaneously targets cosmological parameters and the initial matter distribution.
@@ 2.1 Velocities in Lagrangian Perturbation Theory
Note that the growth factor $D$ as well as the growth rate $f$ are obtained from the $\Lambda$CDM prediction. The EdS approximation has been shown to be accurate to the level of $0.1-0.2\%$ (at $z=0$) in predicting the power spectrum [34], while significantly simplifying the numerical computations. From here on, we always assume a fixed time or redshift at which the forward model is evaluated and hence drop the explicit time argument.
@@ 2.2 Accuracy of the Eulerian-frame momentum
To evaluate the LPT implementation and its convergence, we compare it to a full N-body simulation run for the same volume and cosmology with an identical cut-off at $\Lambda$ applied to its initial conditions as in the forward model (see ref. [35] for a detailed description of the reference simulations). The fiducial Euclidean $\Lambda$CDM cosmology assumed throughout this work is identical to refs. [41, 35] and defined by
@@
The comparison between forward model and reference simulation is shown in figure 1 at $z=0.5$, $\Lambda=0.20\,h/\mathrm{Mpc}$, and it is extended to a smaller cut-off, $\Lambda=0.10\,h/\mathrm{Mpc}$, or a higher redshift, $z=1.0$, in the figures 22 and 23, respectively. As expected, the residual power spectrum (top right) decreases as the order of the LPT solution increases. Moreover, we observe the same alternating behavior for the cross-correlation between residuals and the ground truth (lower right) that we already derived for the rest-frame density [35]. It originates from the fact that the residuals at order $n_{\mathrm{LPT}}$ are dominated by $\delta^{(n_{\mathrm{LPT}}+1)}$, and hence their cross-correlation with the ground truth scales at least as $\left\langle\delta^{(2)}\delta^{(2)}\right\rangle\propto k^{4}$ for odd $n_{\mathrm{LPT}}$ and at least as $\left\langle\delta^{(1)}\delta^{(3)}\right\rangle\propto k^{2}P(k)$ for even ones. At $z=0.5$ and $\Lambda=0.20\,h/\mathrm{Mpc}$, the momentum power spectrum is predicted to an accuracy better than 8% at third and better than 4% at fourth order LPT, and the cross-correlation coefficient is better than 98.5%. This means the momentum is somewhat less accurate, by a factor $\sim 4$ in the power spectrum, than the density at identical LPT order. This trend continues at higher redshifts. It can be explained by the prefactor $n$ in eq. (2.4), which enhances the next-order contribution that is no longer included in an $n$-th order forward model by a factor $(n+1)$ relative to the case of the density. The accuracy of the momentum power spectrum improves considerably as one moves to lower cut-offs, at $\Lambda=0.10\,h/\mathrm{Mpc}$ and $z=0.5$ it is at 2% and 1% for the 3LPT and the 4LPT forward model, respectively. Transverse displacement contributions, as figure 1 illustrates, have a negligible impact.
@@ 2.3 Accuracy of the velocity
Isolated, sharp features in real space impact the power spectrum on all scales in an uncontrollable manner. We therefore focus on the Delaunay tessellation exclusively when quantitatively evaluating the accuracy at which the forward model predicts the velocity. In figure 3, we compare the LOS velocity between the forward model and the reference simulations at $z=0.5$ for a cut-off $\Lambda=0.20\,h/\mathrm{Mpc}$. In figures 22 and 23 we extend the comparison to a lower cut-off, $\Lambda=0.10\,h/\mathrm{Mpc}$ and a higher redshift, $z=1.0$, respectively. With increasing LPT order the magnitude of the residual power spectrum decreases, as expected. The magnitude of the residuals relative to the ground truth is similar for momentum and velocity. They also scale similarly at low wavenumbers; at high wavenumbers the relative residuals increase slightly less steeply for the velocities compared to momentum (figure 1) and density (not shown here). This is in particular the case for $n_{\mathrm{LPT}}\geq 3$. The power spectrum difference of the velocities oscillates around zero between LPT orders, in particular on intermediate scales. That is, for $n_{\mathrm{LPT}}=1,3$ the velocity power spectrum is over-predicted and under-predicted for $n_{\mathrm{LPT}}=2,4$. In contrast, the density and momentum power spectra are always under-predicted. However, in both cases, the power spectrum differences approach zero from below or above for $n_{\mathrm{LPT}}$ being even or odd, respectively. Interestingly, the velocity power spectrum is predicted quite accurately at 1LPT order, but the residuals are huge and decrease by more than an order of magnitude when proceeding to second order. This indicates that the good match of $P_{u_{\parallel}}$ at 1LPT is just a coincidence. For the 4LPT model at $z=0.5$, $\Lambda=0.20\,h/\mathrm{Mpc}$ the accuracy for the velocity power spectrum is better than $3\%$ and it improves to percent-level accuracy for $\Lambda=0.10\,h/\mathrm{Mpc}$ or $z=1.0$.
@@ 2.4 Scaling of the leading-order counter terms
$\left\langle\epsilon_{\alpha}\left(\bm{\bm{k}}\right)\epsilon_{\alpha}\left(\bm{\bm{k}}^{\prime}\right)\right\rangle=\left(2\pi\right)^{3}\delta^{\mathrm{D}}\left(\bm{\bm{k}}-\bm{\bm{k^{\prime}}}\right)P_{\epsilon_{\alpha}}(k)\propto\begin{cases}k^{2}&\quad\mathrm{for}\quad\alpha=\pi\\ k^{4}&\quad\mathrm{for}\quad\alpha=\delta\,,\end{cases}$ (2.10)
@@ 3.1 Summary of the redshift-space model
$1+\tilde{\delta}_{\mathrm{g},\mathrm{det}}\left(\tilde{\bm{\bm{x}}}\right)$ $=\int d^{3}\bm{\bm{q}}~{}\delta^{\mathrm{D}}\left[\tilde{\bm{\bm{x}}}-\bm{\bm{q}}-\bm{\bm{s}}\left(\bm{\bm{q}}\right)-u_{\mathrm{g},\mathrm{det}\parallel}\left(\bm{\bm{q}}\right)\hat{\bm{\bm{n}}}\right]\,\left[1+\delta_{\mathrm{g},\mathrm{det}}\left(\bm{\bm{q}}\right)\right]$
@@
The Jacobian factor of the displacement to redshift space $\left(1+\partial_{x_{\parallel}}u_{\mathrm{g},\mathrm{det}\parallel}\right)^{-1}$ ensures the correct redshift-space density. As in the rest-frame implementation of Lagrangian bias, the bias operators are multiplied by an additional Jacobian factor $1+\delta_{\mathrm{m}}$ from the Lagrangian-to-rest frame mapping. At leading order, they correspond to the desired Lagrangian operators shifted to redshift space $\mathcal{O}\left[\bm{\bm{q}}\left(\tilde{\bm{\bm{x}}}\right)\right]$, and at higher orders the multiplication by $\delta_{\mathrm{m}}$ can be absorbed by a redefinition of the bias coefficients (note that this prefactor is the rest-frame matter density evaluated at the redshift-space position). In the analysis, all bias coefficients are varied separately, and to implement this efficiently we compute the displacements of all terms in eq. (3.7) independently. In this way, we can scan through different bias coefficients (at fixed cosmology) more efficiently without the need to reevaluate the displacement at every new value tested.
@@
$\delta^{(1)}_{\Lambda}$$\left\{\sigma^{(n)},\bm{\bm{t}}^{(n)},\right.$$\left.\sigma^{\prime(n)},\bm{\bm{t}}^{\prime(n)}\right\}$$\left\{\bm{\bm{s}}^{(n)}\right\}$$\left\{\mathcal{O}^{\mathrm{L}}\right\}$$\left\{u^{(n)}_{\parallel},\mathcal{U}\right\}$$u_{\mathrm{g},\mathrm{det}\parallel}$$+$$\mathbb{1}\,,\left\{\mathcal{O}^{\mathrm{L}}\right\}$$\left\{\mathcal{O}^{\mathrm{L}}\right\}$$\tilde{\delta}_{\mathrm{g},\mathrm{det}}$ $\Lambda$ $N_{\mathrm{G,in}}$ $N_{\mathrm{G,fwd}}$, $n_{\mathrm{LPT}}$ $o_{\mathrm{bias}}$ $o_{u,\mathrm{bias}}$ $N_{\mathrm{G,Eul}}$ $k_{\mathrm{max}}$ $N_{\mathrm{G,LH}}$ Fourier-space top hat filter initial grid, $N_{\mathrm{G},\mathrm{Ny}}\left(\Lambda\right)$ LPT grid, aliasing limit for $3/2\,\Lambda$ LPT order $n_{\mathrm{LPT}}\geq o_{\mathrm{bias}}-1$ Lagrangian bias expansion order velocity bias expansion order displacement resolution, $3/2\,N_{\mathrm{G},\mathrm{Ny}}\left(\Lambda\right)$ density assignment, NUFFT scheme likelihood filter, $k_{\mathrm{max}}=\Lambda$ likelihood grid, $N_{\mathrm{G},\mathrm{Ny}}\left(k_{\mathrm{max}}\right)$ numerical parameter redshift space forward model meaning & best-practice
@@ 3.2 Resolution effects
[…]
@@ 4 Growth-rate inference at fixed initial conditions
The forward model residuals (section 3.3) are indicative of the accuracy of the model, but are not easily translated quantitatively into the accuracy and precision that we can expect in a field-level parameter inference. To this end, we perform a set of parameter recovery tests from halos in redshift space in N-body simulations. The data set and analysis setup are detailed in section 4.1 and we present the results in section 4.2.
@@ 4.1 Analysis
[…]
The analysis of simulated data allows us to fix the initial conditions $\delta^{(1)}$ to the known ground truth. This removes a considerable amount of uncertainty from the final parameter constraints while speeding up the analysis; hence we are able to perform very stringent tests and to scan through a wide range of analysis configurations. Unbiased cosmological constraints at fixed initial conditions are an important prerequisite for the full cosmological analysis, but they do not automatically guarantee consistent results when the initial conditions are varied simultaneously. The simultaneous inference of cosmological parameters and $\delta^{(1)}$ was demonstrated recently in the rest frame [27, 32, 33], and we plan to explore such an analysis in redshift space in the near future.
To derive constraints on the growth rate $f$, we marginalize over the noise amplitudes in eq. (3.11) and over the density and velocity bias coefficients $b_{\mathcal{O}}$ and $\beta_{\mathcal{U}}$. The coefficients $b_{\mathcal{O}}$ enter the likelihood quadratically and can be marginalized analytically [66]. We choose this option as it has shown to reduce the correlation lengths of sampling chains [27]. The noise amplitudes and the coefficients $\beta_{\mathcal{U}}$ are sampled explicitly. We report our results in terms of mode-centered 68% credible intervals; further details on the sampling algorithm, stopping criteria and convergence tests are provided in appendix E.
@@ 4.2 Results
  • All results presented in this section so far neglect the transverse contribution in the 3LPT and the 4LPT forward model, which was found to have a negligible impact on the model accuracy in section 3.3. We verify that this also holds for the growth-rate inference in figure 31.
@@
Further, the dependence of the inferred growth rate on $\Lambda$ at fixed $k_{\mathrm{max}}$ indicates some residual impact of small-scale modes beyond $k_{\mathrm{max}}$. This effect is not absorbed by the higher-order bias terms which we explore in figures 12, 26 and 27. It might point to an issue with the simplified noise model (likelihood) adopted here. Specifically, we neglect noise contributions from the RSD displacement in eq. (3.12). At next-to-leading order, these take the form $(\partial_{x_{\parallel}}u_{\parallel})(\tilde{\bm{\bm{x}}})\,\epsilon(\tilde{\bm{\bm{x}}})$; the term directly correlates with the velocity and depends linearly on the parameter of interest, $f$. It is plausible that this noise term, if unaccounted, impacts the result. We plan to develop a noise model for LEFTfield that can account for scale dependence and the noise properties in redshift space simultaneously. This will allow us to explore the dependency of the remaining percent-level systematics on the noise description. We also plan to investigate the convergence of the forward model and the parameter inference for higher LPT orders.

### v2

@@ Abstract
Forward modeling the galaxy density within the Effective Field Theory of Large Scale Structure (EFT of LSS) enables field-level analyses that are robust to theoretical uncertainties. At the same time, they can maximize the constraining power from galaxy clustering on the scales amenable to perturbation theory. In order to apply the method to galaxy surveys, the forward model must account for the full observational complexity of the data. In this context, a major challenge is the inclusion of redshift space distortions (RSDs) from the peculiar motion of galaxies. Here, we present improvements in the efficiency and accuracy of the RSD modeling in the perturbative LEFTfield forward model. We perform a detailed quantification of the perturbative and numerical error for the prediction of momentum, velocity and the redshift-space matter density. Further, we test the recovery of cosmological parameters at the field level, namely the growth rate $f$, from simulated halos in redshift space. For a rigorous test and to scan through a wide range of analysis choices, we fix the linear (initial) density field to the known ground truth but marginalize over all unknown bias coefficients and noise amplitudes. With a third-order model for gravity and bias, our results yield $1\,\%$ statistical and $1.5\,\%$ systematic error. The computational cost of the redshift-space forward model is only $\sim 1.5$ times of the rest frame equivalent, enabling future field-level inference that simultaneously targets cosmological parameters and the initial matter distribution.
@@ 2.1 Velocities in Lagrangian Perturbation Theory
[…]
@@ 2.2 Accuracy of the Eulerian-frame momentum
[…]
@@
The comparison between forward model and reference simulation is shown in figure 1 at $z=0.5$, $\Lambda=0.20\,h/\mathrm{Mpc}$, and it is extended to a smaller cut-off, $\Lambda=0.10\,h/\mathrm{Mpc}$, or a higher redshift, $z=1.0$, in the figures 23 and 24, respectively. As expected, the residual power spectrum (top right) decreases as the order of the LPT solution increases. Moreover, we observe the same alternating behavior for the cross-correlation between residuals and the ground truth (lower right) that we already derived for the rest-frame density [35]. It originates from the fact that the residuals at order $n_{\mathrm{LPT}}$ are dominated by $\delta^{(n_{\mathrm{LPT}}+1)}$, and hence their cross-correlation with the ground truth scales at least as $\left\langle\delta^{(2)}\delta^{(2)}\right\rangle\propto k^{4}$ for odd $n_{\mathrm{LPT}}$ and at least as $\left\langle\delta^{(1)}\delta^{(3)}\right\rangle\propto k^{2}P(k)$ for even ones. At $z=0.5$ and $\Lambda=0.20\,h/\mathrm{Mpc}$, the momentum power spectrum is predicted to an accuracy better than 8% at third and better than 4% at fourth order LPT, and the cross-correlation coefficient is better than 98.5%. This means the momentum is somewhat less accurate, by a factor $\sim 4$ in the power spectrum, than the density at identical LPT order. This trend continues at higher redshifts. It can be explained by the prefactor $n$ in eq. (2.4), which enhances the next-order contribution that is no longer included in an $n$-th order forward model by a factor $(n+1)$ relative to the case of the density. The accuracy of the momentum power spectrum improves considerably as one moves to lower cut-offs, at $\Lambda=0.10\,h/\mathrm{Mpc}$ and $z=0.5$ it is at 2% and 1% for the 3LPT and the 4LPT forward model, respectively. Transverse displacement contributions, as figure 1 illustrates, have a negligible impact.
@@ 2.3 Accuracy of the velocity
Isolated, sharp features in real space impact the power spectrum on all scales in an uncontrollable manner. We therefore focus on the Delaunay tessellation exclusively when quantitatively evaluating the accuracy at which the forward model predicts the velocity. In figure 3, we compare the LOS velocity between the forward model and the reference simulations at $z=0.5$ for a cut-off $\Lambda=0.20\,h/\mathrm{Mpc}$. In figures 23 and 24 we extend the comparison to a lower cut-off, $\Lambda=0.10\,h/\mathrm{Mpc}$ and a higher redshift, $z=1.0$, respectively. With increasing LPT order the magnitude of the residual power spectrum decreases, as expected. The magnitude of the residuals relative to the ground truth is similar for momentum and velocity. They also scale similarly at low wavenumbers; at high wavenumbers the relative residuals increase slightly less steeply for the velocities compared to momentum (figure 1) and density (not shown here). This is in particular the case for $n_{\mathrm{LPT}}\geq 3$. The power spectrum difference of the velocities oscillates around zero between LPT orders, in particular on intermediate scales. That is, for $n_{\mathrm{LPT}}=1,3$ the velocity power spectrum is over-predicted and under-predicted for $n_{\mathrm{LPT}}=2,4$. In contrast, the density and momentum power spectra are always under-predicted. However, in both cases, the power spectrum differences approach zero from below or above for $n_{\mathrm{LPT}}$ being even or odd, respectively. Interestingly, the velocity power spectrum is predicted quite accurately at 1LPT order, but the residuals are huge and decrease by more than an order of magnitude when proceeding to second order. This indicates that the good match of $P_{u_{\parallel}}$ at 1LPT is just a coincidence. For the 4LPT model at $z=0.5$, $\Lambda=0.20\,h/\mathrm{Mpc}$ the accuracy for the velocity power spectrum is better than $3\%$ and it improves to percent-level accuracy for $\Lambda=0.10\,h/\mathrm{Mpc}$ or $z=1.0$.
@@ 2.4 Scaling of the leading-order counter terms
[…]
@@ 3.1 Summary of the redshift-space model
[…]
@@
[…]
@@
$\delta^{(1)}_{\Lambda}$$\left\{\sigma^{(n)},\bm{\bm{t}}^{(n)},\right$$\left.\sigma^{\prime(n)},\bm{\bm{t}}^{\prime(n)}\right\}$$\left\{\bm{\bm{s}}^{(n)}\right\}$$\left\{\mathcal{O}^{\mathrm{L}}\right\}$$\left\{u^{(n)}_{\parallel},\mathcal{U}\right\}$$u_{\mathrm{g},\mathrm{det}\parallel}$$\mathbb{1}\,,\left\{\mathcal{O}^{\mathrm{L}}\right\}$$\left\{\mathcal{O}^{\mathrm{L}}\right\}$$\tilde{\delta}_{\mathrm{g},\mathrm{det}}$ $\Lambda$ $N_{\mathrm{G,in}}$ $N_{\mathrm{G,fwd}}$, $n_{\mathrm{LPT}}$ $o_{\mathrm{bias}}$ $o_{u,\mathrm{bias}}$ $N_{\mathrm{G,Eul}}$ $k_{\mathrm{max}}$ $N_{\mathrm{G,LH}}$ Fourier-space top hat filter initial grid, $N_{\mathrm{G},\mathrm{Ny}}\left(\Lambda\right)$ LPT grid, aliasing limit for $3/2\,\Lambda$ LPT order $n_{\mathrm{LPT}}\geq o_{\mathrm{bias}}-1$ Lagrangian bias expansion order velocity bias expansion order displacement resolution, $3/2\,N_{\mathrm{G},\mathrm{Ny}}\left(\Lambda\right)$ density assignment, NUFFT scheme likelihood filter, $k_{\mathrm{max}}=\Lambda$ likelihood grid, $N_{\mathrm{G},\mathrm{Ny}}\left(k_{\mathrm{max}}\right)$ numerical parameter redshift space forward model meaning & best-practice
@@ 3.2 Resolution effects
where $L_{\mathrm{box}}$ is the side length of a cubic simulation box and $\lceil\ldots\rceil$ indicates up-rounding to the next integer. One the other hand, if a field has support up to $n\Lambda$ and we are only interested in modes below $\Lambda$, it is sufficient to represent the field on a grid large enough to avoid aliasing of high-$k$ modes into the modes of interest. The corresponding minimum grid size is $N_{\mathrm{G},\mathrm{Ny}}\left(k_{\mathrm{alias}}\right)$ with
@@ 4 Growth-rate inference at fixed initial conditions
The forward model residuals (section 3.3) are indicative of the accuracy of the model, but are not easily translated quantitatively into the accuracy and precision that we can expect in a field-level parameter inference. To this end, we perform a set of parameter recovery tests from halos in redshift space in N-body simulations. The data set and analysis setup are detailed in section 4.1, we present the results in section 4.2 and put the precision of our tests in context in section 4.3.
@@ 4.1 Analysis
We test the forward model by inferring the growth rate $f$ from N-body halos in redshift space. The halos provide a non-trivially biased tracer of the underlying non-linear matter density. We have two simulation realizations available, for which we consider three redshift snapshots at z=0.0, 0.5, 1.0z=0.0,\,0.5,\,1.0. Halos are identified in their rest frame with the ROCKSTAR halo finder [67], in contrast to previous redshift-space analyses with LEFTfield that used the same simulations but focused on AMIGA halos [68]. In particular we consider two halo mass bins, $\lg M/\left(h^{-1}M_{\odot}\right)\in\left[12.5,13.0\right]$ and $\lg M/\left(h^{-1}M_{\odot}\right)\in\left[13.0,13.5\right]$. To boost the halos to redshift space, we displace their centers by the center-of-mass velocity of the halo core, i.e. the mass-averaged velocity within 10% of the halo radius. We provide a detailed comparison with results from the two-step model [41] on this data set in appendix D.2. For further details on the simulations see ref. [35] and appendix F therein.
To derive constraints on the growth rate $f$, we marginalize over the noise amplitudes in eq. (3.11) and over the density and velocity bias coefficients $b_{\mathcal{O}}$ and $\beta_{\mathcal{U}}$. The coefficients $b_{\mathcal{O}}$ enter the likelihood quadratically and can be marginalized analytically [66]. We choose this option as it has shown to reduce the correlation lengths of sampling chains [27]. The noise amplitudes and the coefficients $\beta_{\mathcal{U}}$ are sampled explicitly. We fix the initial conditions $\delta^{(1)}$ to the known ground truth which allows us to perform tests at a high precision while simultaneously scanning through a wide range of analysis configurations, see also section 4.3. We report our results in terms of mode-centered 68% credible intervals; further details on the sampling algorithm, stopping criteria and convergence tests are provided in appendix E.
@@ 4.2 Results
  • All results presented in this section so far neglect the transverse contribution in the 3LPT and the 4LPT forward model, which was found to have a negligible impact on the model accuracy in section $3.$ We verify that this also holds for the growth-rate inference in figure 32.
@@
[…]
4.3 On the precision of the growth rate inference
The analysis of simulated data allows us to fix the initial conditions $\delta^{(1)}$ to the known ground truth. Unbiased cosmological constraints at fixed initial conditions are an important prerequisite for the full cosmological analysis, but they do not automatically guarantee consistent results when the initial conditions are varied simultaneously. The simultaneous inference of cosmological parameters and $\delta^{(1)}$ was demonstrated recently in the rest frame [27, 32, 33], and we plan to explore such an analysis in redshift space in the near future.
[…]
Figure 14: Comparison between field-level and power- and bispectrum constraints at fixed initial conditions. We consider the lower halo mass bin at $z=0.5$, $k_{\mathrm{max}}=0.12\,h/\mathrm{Mpc}$. For the field-level constraints, unlike the rest of the paper, we simultaneously infer the linear density-field amplitude $\sigma_{8}$ and growth rate $f$ while marginalizing over noise amplitudes and the full set of third-order bias parameters. For the power- and bispectrum we perform a Fisher analysis where the covariance has been adjusted to account for fixed initial conditions, and adopt and marginalize over a bias expansion up to second order and only the leading noise term. We include the power spectrum monopole, quadrupole and hexadecapole as well as the bispectrum monopole in the data vector. In both cases, we use the 3LPT forward model with leading-order velocity bias and set $\Lambda=k_{\mathrm{max}}$.
What our tests do not yet answer is the constraining power of field level analysis when applied under realistic conditions, that is inferring initial conditions and cosmology simultaneously from volumes that correspond to current stage IV surveys. So far, the information accessible to field-level analysis versus state-of-the-art power- and bispectrum (P+B) measurements has, in the context of the EFT approach, only been investigated in the restframe [78, 33, 79, 80] (see [39] for related field-level results in redshift space using a heuristic galaxy forward model). While we leave a full investigation of the redshift-space field-level information for future work, we can estimate the constraining power of the P+B analysis for our setup by performing a Fisher forecast. In the forecast, we use a modified covariance that accounts for fixed initial conditions [81]. We further restrict the bias expansion to second order for the P+B analysis, see section C.1, including the next-to-leading order derivative operator $\nabla_{q}^{2}\sigma$, the leading velocity operator (eq. C.1) and the leading white noise term from eq. (3.11). Their fiducial values are set to match the lower halo mass bin at $z=0.5$ and obtained by a field-level measurement at fixed cosmology. For the field-level results we consider the full third-order bias expansion with scale- and direction-dependent noise terms. [^2] Specifically, we consider the 3LPT results of figure 10 in this comparison. Focusing on the lower mass bin at $k_{\mathrm{max}}=0.12\,h/\mathrm{Mpc}$, $z=0.5$, we obtain, at fixed $\sigma_{8}$, $\sigma_{\mathrm{P+B}}(f)/f_{\mathrm{fid}}=1.1\%$ in comparison to the $0.6\%$ measurement by the field-level analysis.
Notice that up to this point and throughout this paper, we have kept the amplitude of initial fluctuations $\sigma_{8}$ fixed, as our goal was to perform a rigorous test of the redshift-space modeling. The joint inference of $\sigma_{8}$ and $f$, corresponding in a sense to a consistency test of GR, yields even more substantial gains from the field-level analysis over power spectrum and bispectrum (figure 14). This gain is obtained due to the effective breaking of the linear-order degeneracy between the two parameters present for biased tracers. We emphasize again that these results, while a fair comparison, are strictly for the case of fixed initial conditions.